Question 212.56 pts

On January 8, 2018, Sam, age 62, sold for $410,000 his personal residence which had an adjusted basis of $150,000. Sam purchased the home in 2013 and used it as his personal principal residence for the last three years. On May 1, 2018, he purchased a new residence for $520,000. For 2018, Sam should recognize a gain on the sale of his residence of:

$10,000

$250,000

$260,000

$0

Flag this Question

Question 222.56 pts

Jody purchased a home in a suburb of Milwaukee for $300,000 on October 1, 2016. Eighteen months later, her employer transferred her to Minneapolis and Jody sold her home for $340,000. How much of Jody's gain form the sale of her Milwaukee home can be excluded from her gross income?

$0

$10,000

$30,000

$40,000

Flag this Question

Question 232.56 pts

Bobby and Betty Bennett sold for $450,000 in October of 2018 their residence that they had purchased in 2008 for $200,000. They made major capital improvements during their 10-year ownership totaling $40,000. What is their recognized gain?

$0

$210,000

$250,000

$450,000

None of the above.

Flag this Question

Question 242.56 pts

Suppose instead that in the preceding problem the Bennetts sold their home for $800,000. They moved into a smaller home costing $250,000. How much gain must they recognize?

$560,000

$60,000

$500,000

$310,000

None of the above.

Flag this Question

Question 252.56 pts

Assume instead that the Bennetts resided in a very depressed neighborhood and the home purchased in 2008 for $200,000 (capital improvements of $40,000) was sold for only $110,000. How much loss is recognized?

$200,000

$130,000

$0

$90,000

None of the above.
Answer #1

First question

Recognzied gain by sam = 410,000 – 150,000 – 250,000 = 10,000

Second question:

Recognized gain = $0

Third:

Recognized gain by Bobby and Betty = 450,000 - 40,000 – 200,000 – 500,000 (exclusion) = $0

Fourth:

Recognized gain by Bobby and Betty = 800,000 - 40,000 – 200,000 – 500,000 (exclusion) = $60,000

Fifth:

Loss recognized = $0